A certain ion X3+ contains 24 electrons and 32 neutrons. The chemical symbol and the mass number of this ion are:


1
Expert's answer
2022-06-13T11:03:04-0400

Solution:

The net charge of an ion is given by the difference that exists between the number of protons located inside its nucleus and the number of electrons surrounding its nucleus.

Therefore,

net charge = number of protons − number of electrons

Therefore,

(+3) = number of protons − 24

number of protons = 24 + 3 = 27


Thus, the element X has 27 protons inside its nucleus, i.e. an atomic number equal to 27.

The element with atomic number 27 is cobalt (Co)

Thus, X3+ ion is Co3+ ion


The mass number of an atom is given by the number of nucleons (protons and neutrons) located inside its nucleus.

Therefore,

mass number = number of protons + number of neutrons

mass number = 27 + 32 = 59

mass number = 59


Answer:

The chemical symbol of this ion is Co3+

The mass number of this ion is 59